OpenStudy (anonymous):

Which of the following polynomials has a degree of 3? A. x3y3 - 3xy B. x + y + z C. x2y + 5y2 D. x2y2 - y3

hero (hero):

@tm-cool, which do you believe is correct?

OpenStudy (anonymous):

i think a

hero (hero):

No, the exponents of the first term have to add to 3 in order for the degree of polynomial to be three. For a you have an exponent of 3 for x and an exponent of 3 for y. 3 + 3 = 6 so, a can't be it.

OpenStudy (anonymous):

so the its d

OpenStudy (shamil98):

Would it not be B? x + y + z

hero (hero):

No, it's not d because 2 + 2 = 4

hero (hero):

@tm-cool, you should have gotten it on the second try. I gave you enough of a hint.

OpenStudy (anonymous):

i got it i didnt your answer right

hero (hero):

I haven't mentioned what the answer is. Which answer choice do you believe is correct?

OpenStudy (anonymous):

b

hero (hero):

It's not b either. And by process of elimination you know that c is the answer. But guessing is not the way.

OpenStudy (anonymous):

ok none of them equaled three c 2+2=4

hero (hero):

Ugh, this problem isn't what I thought it was... d is correct. I didn't see the cube term

hero (hero):

c would be correct if d didn't have the cubed second term.

hero (hero):

Sorry for mis-leading you

OpenStudy (anonymous):

ok

hero (hero):

Here's the proof that c also has degree of 3: http://www.wolframalpha.com/input/?i=degree+of+x2y+%2B+5y2

hero (hero):

So those answer choices are misleading. c and d both have degree of 3

hero (hero):

By the way, any time I say "first term" in reference to polynomials, I mean to say "leading term".

hero (hero):

Here's more info about it on wikipedia: http://en.wikipedia.org/wiki/Degree_of_a_polynomial

hero (hero):

Also if you're not going to use \(\LaTeX\) to post your questions, you should use the caret symbol to emphasize the exponents. x3y3 is not a proper notation for \(x^3y^3\). However x^3y^3 is a proper notation.